# Let’s Get to the Point: LLM-Supported Planning, Drafting, and Revising of Research-Paper Blog Posts

### v1

@@ 6.4.2. RQ2: Real-World Tool Perceptions
Overall, we observed benefits from Papers-to-Posts in fulfilling Papers-to-Posts’s design goals. Regarding the first design goal (DG1), participants indicated only a slight improvement in feeling confident that all essential information was covered with Papers-to-Posts (M=5.00,Q1=4.50,Q3=6.00) as opposed to the baseline (M=5.00,Q1=4.25,Q3=6.00). Given that the automatic method for parsing the paper PDF in the deployment study was likely to make more errors than the partially Wizard-of-Oz method for parsing papers in the lab study, it makes sense that participants would now see a reduced benefit from Papers-to-Posts’s paper outline in terms of ensuring all relevant information is included. Multiple participants under both conditions noted that the parsing of their papers was incomplete or flawed. However, participants still indicated an increased ease in incorporating information from the paper in the blog post with Papers-to-Posts (M=6.00,Q1=5.00,Q3=6.00) compared to the baseline (M=4.50,Q1=3.25,Q3=5.75). For instance, P14 commented, ”I liked being able to check the boxes of the portions of the paper I wanted to include in each section.”
With respect to our second design goal (DG2), there was no difference in the median for how well Papers-to-Posts (M=5.00,Q1=4.00, Q3=6.00) versus the baseline (M=5.00,Q1=5.00,Q3=6.00) facilitated production of appropriate text, unlike in the lab study (Figure 8e-h). Nevertheless, in their qualitative feedback, several participants noted how Papers-to-Posts’s generation and modification features helped them to create more appropriate text. For instance, P24 explained that the ”preset custom instructions was very useful because can need to be very specific using these tools,” and P14 noted that the ”provided buttons for making text more or less dramatic was nice.” Furthermore, in their survey responses, participants indicated that Papers-to-Posts (M=6.00,Q1=4.50,Q3=6.00) facilitated iterating on the blog post more so than the baseline (M=5.00,Q1=3.50,Q3=6.00), which is reflected in the lower number of writing actions participants took for a unit change in their initial drafts that we observed in Figure 9(c).
7. Discussion and Design Implications
@@
Based on the results of the lab and deployment studies, we reflect on design implications for other human-LLM scientific long-document summarization tools with long-form outputs.
@@ 7.1. Post-Editing is not Enough
Both Papers-to-Posts and the baseline tool included initial LLM-generated drafts generated with powerful LLMs, GPT-4 and GPT-4-32k respectively. Thus, the existence of an initial LLM draft cannot be the primary reason why participants preferred the blog post generated with Papers-to-Posts as compared to the baseline tool. While the participants had the option to simply post-edit the initial draft and refrain from interacting further with the LLM, only two baseline and two treatment participants across the lab and deployment studies did so. This suggests that additional user interfaces and interaction designs can provide better scaffolding supports to further improve upon LLM-generated text compared to simple post-editing.
7.2. Long-Document Outlines
While blog post writing has additional goals and constraints compared to general document summarization, our work also has implications that may be useful for the task of general human-AI long document summarization. Specifically, a main challenge here is the high costs of human summarizer reviewing, locating, and utilizing different essential information in a long document. Prior work in human-LLM summarization has attempted to address this challenge in various manners, including presenting visualizations for the document to be summarized (Zhang et al., 2023b, 2011), dividing the document into chunks (Cheng et al., 2023), and providing document highlights for information already included in the summary (Norkute et al., 2021). However, to the best of our knowledge, no prior work has explored the use of generated outlines to help users review the document to be summarized.
In this work, participants found it easier to find and incorporate essential information from the document using Papers-to-Posts and participants consistently noted the selectable bullet points as one of the most useful features. We found that, given this tool, participants were able to make more change to their initial drafts for a given amount of time or writing actions. We therefore note that presenting LLM-generated outlines of scientific long documents may make the planning step of writing their summaries easier. As participants faced confusion about the connection between the outline and the paper text, we also recommend working to make this connection as clear as possible. How detailed the outline should be remains uncertain. While several participants appreciated the bullet points, a couple mentioned that the outline was overwhelming. As one participant proposed, a hierarchy of bullet points may be more useful.
7.3. Transparent Affordances for Long-form Writing Sub-Tasks
In contrast with shorter forms of scientific writing such as tweetorials (Long et al., 2023) and extended metaphors (Kim et al., 2023) explored in prior work on human-LLM scientific writing, longer scientific texts such as research-paper blog posts require completing a number of sub-tasks common to long-form science communication. For example, because research-paper blog posts are often for broader audiences than the paper itself, Papers-to-Posts includes a ”simpler terms” option to make large swaths of text more digestible for such readers. As another example, because different sections of a research-paper blog post may serve different purposes, we pre-populated the custom instructions for each of the initial draft sections (e.g., for the introduction section, we ask the tool to present a hook, a high-level description of the problem and why it is interesting, etc.). These transparent, section-specific instructions could then be verified or adjusted by the user. Several participants commented on the usefulness of Papers-to-Posts’s different affordances for streamlining different generation and modification sub-tasks such as these. Thus, we observe that similar tools may assist users by including affordances for sub-tasks related to generating and modifying long-form output.
@@ 8. Limitations and Future Work
The majority of the participants were PhD students in computer science. Although many of the takeaways discussed in this paper may generalize to other groups of researchers, it would be interesting to see how more diverse populations surface unique needs from the tool. For example, researchers in the field of computer science tend to be better versed in artificial intelligence than those in other fields, so they may not need as much guidance in utilizing an LLM writing tool. Different areas of research may also have very different traditional paper and blog post structures compared to computer science, which may benefit from different tool affordances. Furthermore, PhD students are prime candidates for writing blog posts, which can help them raise awareness of their work as they are starting their research careers. However, more senior researchers may desire specific features for writing their blog posts, such as a feature to coordinate discussion of multiple works or a feature to connect an older work to the current research landscape. We also leave it to future work to investigate how different aspects of the input paper such as length, recency, and author position may affect the support that users want from the tool.
@@ 9. Ethical Considerations
There are a few ethical considerations that we would like to note regarding this work. For one, it is well-known that LLMs may hallucinate information included in their generated text (Bubeck et al., 2023). We mitigate this risk by focusing on researchers summarizing their own work, which they should know well; for example, multiple participants noted inaccuracies or exaggerations in generated text, but most were able to correct those issues. A couple participants raised concerns that they had not caught all the errors by the end of their session, but as one of them noted, with more time they could be thorough in ensuring that the text is correct. We believe that LLM usage will increase in the future, and it is important to study effective design. Another ethical concern is whether using an LLM tool to write research-paper blog posts could lead to the loss of important details from the paper. P3 in the baseline condition raised this point: ”…it seems like the easier you make it, i.e., the better the initial summary, the less likely I am to myself remember those things that happen to be missing.” That said, participants overall felt more confident that all essential information was covered in their blog posts with Papers-to-Posts, so this work provided one step in the direction of support for reducing missing information. Furthermore, an LLM tool for writing research-paper blog posts may contribute to making science more accessible in society. This may raise concerns of reducing paper readership. However, blog posts do not provide the same amount of information as papers and do not go through peer review, research papers still hold their own value, and blog posts can act as catalysts for paper reading. Finally, researchers may not acknowledge the use of AI in writing their blog posts, which could mislead readers, particularly if the researcher has not verified the blog post contents. For the participants who shared their blog posts, we found that many chose to disclose its use, but to ensure better transparency around its use, future versions of Papers-to-Posts may highlight the provenance of the output writing (e.g., by visualizing the writer’s interaction with the LLM-powered tool (Naimul Hoque et al., 2023)).
@@ 10. Conclusion
We introduce Papers-to-Posts, an LLM-powered research-paper blog post writing tool, which effectively helps researchers plan, draft, and revise their blog posts, Through a formative study (N=6) with an LLM technology probe for this task, we identified the following challenges faced by writers: they 1) appreciated a paper outline for finding and selecting relevant information from the paper but faced difficulties in navigating it and 2) tried to provide recurrent instructions for different kinds of sub-tasks to the LLM without specific affordances to support these sub-tasks. We formulated design goals to address each of these challenges, namely to help writers 1) plan their writing by facilitating review of the long document’s contents and selection of essential information to include in the writing and 2) draft and revise their writing by streamlining commonly desired instructions to the LLM for generating and modifying text. We validated that our LLM-powered tool Papers-to-Posts addresses these goals and helps authors write blog posts for their own papers through two user studies comparing Papers-to-Posts against a baseline providing an LLM blog post draft and free-form LLM prompting. In a within-subjects, time-controlled lab study (N=20 participants), we found that Papers-to-Posts significantly increased writer satisfaction with the output blog post, without increasing task completion time or cognitive load. We also observe that Papers-to-Posts led to significantly greater change to the initial LLM draft blog post, indicating that this increased satisfaction may be due to the tool facilitating iteration on the blog post. A field deployment study (N=37 blog posts, 26 participants) showed that in a more realistic setting Papers-to-Posts again resulted in greater change in the blog post for a given amount of editing time or number of editing actions. We hope that this work may inform the design of future human-LLM tools for scientific long-document summarization into long-form output.
@@ Acknowledgements.
- Ghazvininejad et al. (2017) Marjan Ghazvininejad, Xing Shi, Jay Priyadarshi, and Kevin Knight. 2017. Hafez: an interactive poetry generation system. In Proceedings of ACL 2017, System Demonstrations. 43–48.
@@
- Google ([n. d.]) Google. [n. d.]. BARD. https://bard.google.com/chat.
@@
- Medium ([n. d.]) Medium. [n. d.]. Medium. https://blog.medium.com/best-practices-for-writing-on-medium-386506ae62b9.
@@
- Norkute et al. (2021) Milda Norkute, Nadja Herger, Leszek Michalak, Andrew Mulder, and Sally Gao. 2021. Towards explainable AI: Assessing the usefulness and impact of added explainability features in legal document summarization. In Extended Abstracts of the 2021 CHI Conference on Human Factors in Computing Systems. 1–7.
@@
- OpenAI ([n. d.]) OpenAI. [n. d.]. chatGPT. https://chat.openai.com/.
@@
- Van Someren et al. (1994) Maarten Van Someren, Yvonne F Barnard, and J Sandberg. 1994. The think aloud method: a practical approach to modelling cognitive. London: AcademicPress 11, 6 (1994).
@@
Appendix A Prompts
A.1. Prompt for Initial LLM Draft for Baseline Tool
Figure 10 shows the prompt for the initial LLM draft for the baseline tool.
A.2. Prompts for Initial LLM Draft for Papers-to-Posts
A.3. Prompts for Working with Papers-to-Posts
Appendix B Additional Log Analysis Plots
Figures 25 and 26 show additional log analysis plots. Figure 27 shows the distribution of behavioral actions taken.
Appendix C Formative Study Technology Probe
Figure 28 shows the formative study technology probe.

### v2

@@ 5.6.2. Deployment Study - Exploratory
Looking at the median results, participants did not indicate an improvement in feeling confident that all essential information was covered with Papers-to-Posts (M=5.00, Q1=4.50, Q3=6.00) as opposed to the baseline (M=5.00, Q1=4.25, Q3=6.00). Given that the automatic method for parsing the paper PDF in the deployment study was likely to make more errors than the partially manual method for parsing papers in the lab study, it makes sense that participants would now see a reduced benefit from Papers-to-Posts’s paper outline in terms of ensuring all relevant information is included. Multiple participants under both conditions noted that the parsing of their papers was incomplete or flawed. However, participants still indicated an increased ease in incorporating information from the paper in the blog post with Papers-to-Posts (M=6.00, Q1=5.00, Q3=6.00) compared to the baseline (M=4.50, Q1=3.25, Q3=5.75). For instance, P14 commented, “I liked being able to check the boxes of the portions of the paper I wanted to include in each section.”
5.6.3. Summary
In both studies, participants indicated increased ease in incorporating paper content in the blog post with Papers-to-Posts. However, the two studies had different outcomes with respect to participants’ confidence in all essential information being covered in the blog post. While lab study participants indicated higher confidence with Papers-to-Posts the deployment study participants did not. This may be because, in the deployment study, we did not include manual corrections to the paper parsing.
5.7. Fulfillment of Drafting and Revising Design Goals (DG2, DG3)
We analyzed participants’ Likert-type responses to questions related to fulfillment of our second and third design goals, related to drafting and revising the blog post (Figure 8c-d,g-h).
5.7.1. Lab Study - Exploratory
Participants’ Likert responses regarding DG2 and DG3 indicate that they found it easier to produce appropriate blog post text (treatment: M=6.00, Q1=5.00, Q3=6.00; baseline: M=5.00, Q1=4.00, Q3=6.00) and iterate on the blog post text (treatment: M=6.00, Q1=5.00, Q3=6.00; baseline: M=5.00, Q1=4.00, Q3=6.00) in the treatment condition (Figure 8c-d). Furthermore, the interaction logs for the treatment condition indicate that participants took advantage of DG2-based affordances for generating text and DG3-based affordances for modifying text (Appendix 25). For an average of 6.85 generations per treatment condition, participants often utilized custom instructions (5.70 times), selected paragraphs or bullet points (6.35 times), and included the context of prior blog post sections (6.00 times), though we note that these features were pre-filled or toggled on by default. In addition, they modified text 6.85 times on average, with the “condense” (1.85 times), “custom” (1.80 times), and “simpler terms” (1.40 times) modifications being the most common.
5.7.2. Deployment Study - Exploratory
There was no difference in the median for how well Papers-to-Posts (M=5.00, Q1=4.00, Q3=6.00) versus the baseline (M=5.00, Q1=5.00, Q3=6.00) facilitated production of appropriate text, unlike in the lab study (Figure 8e-h). Nevertheless, participants indicated that Papers-to-Posts (M=6.00, Q1=4.50, Q3=6.00) facilitated iterating on the blog post more so than the baseline (M=5.00, Q1=3.50, Q3=6.00).
5.7.3. Summary
In both studies, participants using Papers-to-Posts found it easier to iterate on the blog post. However, while lab study participants thought Papers-to-Posts improved their ability to produce appropriate text for the blog post, deployment study participants did not. Perhaps without time constraints, deployment study participants had more time to determine how to prompt the baseline tool or adjust its text in order to produce appropriate text.
5.8. Qualitative Analysis
Across the two studies, participants’ survey and interview comments reflected the difference in editing power between the two tools.
5.8.1. Baseline
Overall, participants using the baseline most commonly cited the initial draft (lab: 6/20, deployment: 8/18) as a helpful aspect of the tool. Thus, the most valued baseline feature was not helpful for editing power. With respect to the initial draft, multiple participants noted editing or adding to the initial draft rather than replacing it. For example, P20-dep explained, “It’s easier to critique than it is to generate,” and P26-dep shared, “It was very useful to have a straw-person of what the blog post could look like (generated by the LLM). I mostly used the draft as a place to edit and steer in the direction I wanted the post to take.” A number of participants also noted the tool’s flexibility as a benefit (lab: 6/20, deployment: 4/18). For example, P18-lab shared, “I was able to provide instructions that fit my mental model of constructing a blog post and it being there as a ‘copilot’ is just what I need.” Flexibility in drafting text may help writers obtain the edits they want. However, the writer holds more responsibility in steering the LLM appropriately; P7-lab reflected, “As a CS student, I like this way more [than the treatment’s way of supporting writing], but I think it needs more thinking to use it creatively.”
5.8.2. Papers-to-Posts’ Planning Features
Participants often noted the selectable paper bullet points and paragraphs as a helpful feature of Papers-to-Posts (lab: 10/20, deployment: 14/19). Several described the outline as helpful for selecting relevant content to generate text. For instance, P15-lab noted, “I could rewrite entire sections or direct it with just a few bullets.” Seven participants mentioned that the outline was useful even just for reviewing the paper’s content or structure. P4-lab commented, “The outline was really helpful in providing a quick overview of what content I had written.” For two lab participants, the bullet points provided transparency around the origins of the initial LLM draft and generations. P12-lab explained, “I liked the bullet point features to see what was being fed in vs not.” By facilitating review and selection of content to include in the blog post, the paper outline may have contributed to participants’ increased editing power.
That said, participants also specified difficulties related to the paper outline. For instance, two lab participants were overwhelmed with the number of bullet points. P9-lab described the outline as “really bloated and hard to navigate.” Meanwhile, three lab participants were confused or encumbered by aspects of utilizing the bullet points. For example, P3-lab “felt stressed by choosing between the entire paragraph at hand and the (presumably) AI-generated summary bullets.”
5.8.3. Papers-to-Posts’ Drafting and Revising Features
Participants frequently mentioned the utility of Papers-to-Posts’s text modification abilities (lab: 12/20, deployment: 9/19). For example, P18-dep reflected, “The ‘MODIFY WRITING’ section is helpful. I like that the interface contains this particular section to support iterating on existing text.” All of the preset modification buttons (“simpler terms”, “condense”, “expand”, “more dramatic”, “less dramatic”) were specifically mentioned as helpful affordances at least once in the lab study alone. The most commonly noted one was the “condense” button. For example, P12-lab commented, “I sometimes put the modify text through multiple rounds of expand and contract, focusing on different parts. I found that helpful.” Three participants commented that they desired more preset modification buttons. For instance, P16-lab wanted a “more technical” button, while P14-dep wanted “more or less casual” options. P7-lab then added, “Or even let user to add their favorite rewrite button.” Interestingly, P2-lab found the preset modification buttons useful for understanding the tool’s capabilities: “It was nice to know what options the tool already knew of, and would perhaps have success with. For example, I would probably not have gone to a blank-box tool and said, ’please make the following more dramatic’. I would have stuck to more basic things like ‘please summarize the following’.” Several participants (lab: 5/20, deployment: 10/19) also noted the usefulness of at least one aspect of the customizable instructions for either generating or modifying text using Papers-to-Posts. P9-shared, “…being able to include explicit instructions for the model to generate text from was helpful in being able to control the information in the text that it generated.” Thus, Papers-to-Posts’s affordances for text drafting and revision may have contributed to participants’ increased editing power.
Participants also discussed difficulties that they encountered in drafting and revising writing with Papers-to-Posts. Four lab participants expressed wanting to take specific actions that were not explicitly supported by Papers-to-Posts, contrasting with the baseline’s flexibility that participants appreciated. As an example, two wanted assistance with checking for redundancies in the blog post. The other two wanted feedback on the blog post writing. Regarding the affordances for modifying text, 8/20 lab and 5/19 deployment participants noted a difficulty. Four lab participants raised confusion about how or when to use certain affordances. As an example, P6-lab was unsure of “when to use which one of the six [modification] options.” Another recurrent issue noted by five lab participants was a modification affordance not working as desired. For P20-lab, setting the desired length of the modified writing to one paragraph still resulted in three, as the back-end prompts are not guaranteed to work as planned. For P12-lab, modifications were sometimes too exaggerated, while for P16-lab, a custom modification to make writing more technical was too subdued. A couple deployment participants also noted that modifying text sometimes introduced errors. Future work may address these points of restriction and confusion noted by the participants using Papers-to-Posts.
6. Discussion
@@
Based on our results, we reflect on design implications for other mixed-initiative tools for research-paper summarization and long-form scientific communication.
@@ 6.1. Post-Editing is Not Enough
Both Papers-to-Posts and the baseline tool included initial LLM-generated drafts generated with powerful LLMs. In the lab study, they were GPT-4 and GPT-4-32k respectively; in the deployment study, both conditions used GPT-4-0125-preview. Nevertheless, while the participants had the option to simply post-edit the initial draft and refrain from interacting further with the LLM, only two baseline and two treatment participants across the lab and deployment studies did so. This suggests that writers may benefit from additional support other than an initial draft when writing research-paper blog posts.
6.2. Increasing Editing Power with Plan-Draft-Revise Workflow
With additional features to support planning, drafting, and revising the blog post, Papers-to-Posts increased writers’ editing power (i.e., change in the blog post per minute or writing action). One reason participants may have demonstrated increased editing power is that we provided an interactive research-paper outline to help them plan content to include in their blog post. Indeed, when using Papers-to-Posts versus the baseline tool, the median participant found it easier to incorporate information from the paper in the blog post. In their qualitative responses, participants also consistently noted the interactive paper outline as one of the most useful features. We therefore note that presenting LLM-generated interactive outlines of research papers shows promise in making the planning step of writing their summaries easier. Another reason participants may have demonstrated increased editing power is that we provided pre-filled yet customizable LLM instructions for generating and modifying the blog post text. In alignment, the median participant found it easier to iterate on the blog post while using Papers-to-Posts. Several participants commented on the usefulness of different aspects of Papers-to-Posts’s preset yet customizable LLM instructions for drafting and revising the blog post text. Thus, we observe that preset LLM instructions scaffolded for customization may support users in drafting and revising long-form scientific communication.
Lowering the temporal cost and writing cost of making an edit, Papers-to-Posts facilitates writers’ iteration on their blog posts, which we saw may be helpful for creating a blog post that satisfies them in a time-constrained setting like the lab study. Iteration may also help the writers to reflect on the blog post text (Zhou and Sterman, 2024). On the other hand, participants often described the baseline as more flexible and allowing them more freedom. Thus, utilizing the two tools in combination may prove useful. Papers-to-Posts may be used when the writer desires suggestions for how to edit their blog post, and the baseline tool (or no LLM tool) could be used when the writer has a specific edit in mind or wants to give themselves creative space to contemplate their own edits (Buschek, 2024).
@@ 7. Limitations and Future Work
Most participants were PhD students in computer science, who may benefit from raising awareness of their early-career work through blog posts; future work may investigate how more senior researchers utilize a tool like Papers-to-Posts. Future studies can also examine how researchers in other fields, who are generally less familiar with LLMs and may have different traditional paper formats, may benefit from a tool like Papers-to-Posts. Furthermore, Papers-to-Posts was designed to support blog posts that summarize a paper, but there are a variety of scientific blog post formats that future work could support (e.g., narrative or opinion-based formats). Future work may additionally investigate how different aspects of the input paper such as length, recency, and author position affect the user experience and outcome.
Papers-to-Posts itself has several limitations. For one, multiple deployment study participants commented on how the automatic paper parsing makes mistakes and misses parts of the paper text. The tool also does not support incorporating figures and tables in the blog post. In addition, future research could work to make the connection between the outline and paper text clearer (e.g., highlight blog post text to reveal related paper text). Moreover, how detailed the paper outline should be remains uncertain. While several participants appreciated the bullet points, a couple mentioned that the outline was overwhelming. Lastly, future work may look into how writers’ expertise may be utilized in supporting their blog post writing; a tool for Q-and-A research-paper summarization, for example, utilizes the researcher’s background to personalize its output (Lim et al., 2024).
@@ 8. Ethical Considerations
It is well-known that LLMs may hallucinate information included in their generated text (Bubeck et al., 2023). We mitigate this risk by focusing on researchers summarizing their own work, which they should know well. Multiple participants identified inaccuracies or exaggerations in generated text, which they could then correct. However, a few participants raised concerns about catching errors, suggesting a need for better support in detecting them. Another ethical concern is whether using an LLM-powered tool to write paper blog posts could lead to the loss of important details. Reflecting on the baseline, P3-lab raised this point: “…it seems like the easier you make it, i.e., the better the initial summary, the less likely I am to myself remember those things that happen to be missing.” Nevertheless, the median participant found it easier to incorporate information from the paper with Papers-to-Posts, so this work provided one step in the direction of reducing loss of important details. Furthermore, while an LLM-powered tool for writing research-paper blog posts may contribute to making science more accessible, it may raise concerns about decreasing paper readership. However, blog posts neither provide the same amount of information nor go through peer review. Papers still hold their own value, and blog posts can act as catalysts for paper reading. Finally, researchers may not acknowledge the use of AI in writing their blog posts, which could mislead readers, particularly if the writer has not properly verified the content. Many participants who shared their blog posts chose to disclose its use, but to provide better transparency, future versions of Papers-to-Posts may highlight the provenance of the output writing (e.g., by visualizing the writer’s interaction with the tool (Naimul Hoque et al., 2023)).
@@ 9. Conclusion
We introduce Papers-to-Posts, an LLM-powered tool for writing research-paper blog posts that follows the Plan-Draft-Revise workflow. We validated that Papers-to-Posts supports these steps through two user studies comparing how researchers write about their own papers with Papers-to-Posts versus a strong baseline (LLM blog post draft and free-form LLM prompting). In a within-subjects lab study (N=20 participants), Papers-to-Posts significantly increased writers’ editing power (i.e., change in initial blog post within a given amount of time), significantly increased satisfaction with the output blog post, and did not increase cognitive load. In a more realistic setting, a deployment study (N=37 blog posts, 26 participants) also showed that Papers-to-Posts resulted in greater editing power in terms of both time and writing actions, without decreasing blog post satisfaction or increasing cognitive load. We hope that this work informs future research and tools related to research-paper summarization and long-form scientific communication.
@@ Acknowledgements.
- Buschek (2024) Daniel Buschek. 2024. Collage is the New Writing: Exploring the Fragmentation of Text and User Interfaces in AI Tools. In Proceedings of the 2024 ACM Designing Interactive Systems Conference. 2719–2737.
@@
- Dhillon et al. (2024) Paramveer S Dhillon, Somayeh Molaei, Jiaqi Li, Maximilian Golub, Shaochun Zheng, and Lionel Peter Robert. 2024. Shaping Human-AI Collaboration: Varied Scaffolding Levels in Co-writing with Language Models. In Proceedings of the CHI Conference on Human Factors in Computing Systems. 1–18.
@@
- Google ([n. d.]) Google. [n. d.]. Gemini. https://gemini.google.com/.
- Google ([n. d.]) Google. [n. d.]. NotebookLM. https://notebooklm.google/.
@@
- Lim et al. (2024) Hyunseung Lim, Ji Yong Cho, Taewan Kim, Jeongeon Park, Hyungyu Shin, Seulgi Choi, Sunghyun Park, Kyungjae Lee, Juho Kim, Moontae Lee, et al. 2024. Co-Creating Question-and-Answer Style Articles with Large Language Models for Research Promotion. In Proceedings of the 2024 ACM Designing Interactive Systems Conference. 975–994.
@@
- Lu et al. (2024) Zhuoran Lu, Sheshera Mysore, Tara Safavi, Jennifer Neville, Longqi Yang, and Mengting Wan. 2024. Corporate Communication Companion (CCC): An LLM-empowered Writing Assistant for Workplace Social Media. arXiv preprint arXiv:2405.04656 (2024).
@@
- Reza et al. (2024) Mohi Reza, Nathan M Laundry, Ilya Musabirov, Peter Dushniku, Zhi Yuan “Michael” Yu, Kashish Mittal, Tovi Grossman, Michael Liut, Anastasia Kuzminykh, and Joseph Jay Williams. 2024. ABScribe: Rapid Exploration & Organization of Multiple Writing Variations in Human-AI Co-Writing Tasks using Large Language Models. In Proceedings of the CHI Conference on Human Factors in Computing Systems. 1–18.
- Shamsabadi and D’Souza (2024) Mahsa Shamsabadi and Jennifer D’Souza. 2024. A fair and free prompt-based research assistant. In International Conference on Applications of Natural Language to Information Systems. Springer, 218–224.
@@
- Zhou and Sterman (2024) David Zhou and Sarah Sterman. 2024. Ai.llude: Investigating Rewriting AI-Generated Text to Support Creative Expression. In Proceedings of the 16th Conference on Creativity & Cognition. 241–254.
@@
Appendix A Exploratory Analysis of Perceived Control and Ownership
We analyzed participants’ Likert-type responses to questions related to control and ownership (Figure 9).
A.1. Lab Study
The median participant did not indicate higher perceived control (treatment: M=5.00, Q1=4.50, Q3=6.00; baseline: M=5.00, Q1=3.00, Q3=6.00) or satisfaction with their control (treatment: M=5.00, Q1=2.75, Q3=6.00; baseline: M=5.00, Q1=3.00, Q3=5.25) under the treatment condition (Figure 9a-b). However, if we look within-subjects, 11 participants had higher perceived control with Papers-to-Posts than the baseline tool (equal: 2, lower: 7), and 10 participants had higher satisfaction with their control using Papers-to-Posts (equal: 4, lower: 6). On the other hand, participants did not demonstrate increased perceived ownership (treatment: M=6.00, Q1=5.75, Q3=6.00; baseline: M=6.00, Q1=5.00, Q3=6.00) or satisfaction with ownership (treatment: M=6.00, Q1=4.00, Q3=6.25; baseline: M=6.00, Q1=5.00, Q3=6.00) under the treatment condition (Figure 9c-d). Although they felt that they had more control over Papers-to-Posts, participants may not necessarily have felt that the tool contributed more or less to the blog post’s creation.
A.2. Deployment Study
As shown in Figure 9e-h, participants indicated higher median satisfaction with control in the treatment condition (treatment: M=5.00, Q1=3.00, Q3=6.00; baseline: M=3.00, Q1=3.00, Q3=4.75), but they showed no difference in median perceived control (treatment: M=5.00, Q1=5.00, Q3=6.00; baseline: M=5.00, Q1=3.00, Q3=5.00), perceived ownership (treatment: M=6.00, Q1=4.50, Q3=6.00; baseline: M=6.00, Q1=6.00, Q3=6.00) or satisfaction with ownership (treatment: M=6.00, Q1=4.00, Q3=6.00; baseline: M=6.00, Q1=5.00, Q3=6.00) in the treatment condition.
A.3. Summary
While lab study participants had higher median perceived control, deployment study participants did not. However, we saw some evidence in both studies of higher perceived satisfaction with control under the treatment condition. Participants may have been more satisfied with their control without perceiving a higher control in the sense that they were more comfortable with letting the tool have a certain amount of control in exchange for more editing power. Across the two studies, we did not see a change in median perceived ownership or satisfaction with ownership.
Appendix B Prompts
B.1. Prompt for Initial LLM Draft for Baseline Tool
B.2. Prompts for Start-Up Step of Papers-to-Posts
B.3. Prompts for Drafting Step of Papers-to-Posts
B.4. Prompts for Revising Step of Papers-to-Posts
Appendix C Additional Log Analysis Plots
Figures 25 and 26 show additional log analysis plots for the lab study, while Figure 27 shows additional log analysis plots for the deployment study.